The length of a rectangle is 3 feet less than twice the width of the rectangle. If the perimeter of the rectangle is 84 feet, find the width and the length.
Width = _____ft
Length=______ft.
x+x+x+x=84
L=2w-3 is what i have come up with but i am not sure where to go from there.

Answer by rapaljer(4671)

Let x = width
2x-3= length

2W + 2L = Perimeter
2(x) + 2(2x-3) = 84
2x+4x-6=84
6x-6=84
6x-6=84+6
6x=90
x=15 ft Width
2x-3 = 2*15-3 = 27 ft Length

Check: 2*15 + 2*27
30 +54
84 It checks!!


For a complete explanation of this and other topics in WORD PROBLEmS, please see my own website. The best way to find it is to use the easy-to-remember and easy-to-spell link www.mathinlivingcolor.com. At the bottom of the page there is a single link that takes you to my Homepage. On my Homepage, look for the link "Basic, Intermediate, and College Algebra: One Step at a Time." Choose "Basic Algebra", and look in "Chapter 1" for "Sections 1.09 and 1.10." Here you will find my own non-traditional explanation that my own students used to find much easier to understand than most of the other textbooks being used. I also have lots of examples, exercises, and answers. In addition, most of the hardest problems are solved in color on the MATH IN LIVING COLOR pages that go with this topic. It's ALL FREE on my website!!!

To contact me please use my Email address at rapaljer@seminolestate.edu.

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus